# Towards Scalable and Stable Parallelization of Nonlinear RNNs

### v1

@@ 4 Scaling and Stabilizing Newton’s Method for Parallel Evaluation
In Section 4.1 we introduce quasi-DEER, a quasi-Newton method that addresses the intractability of DEER for large state sizes. In Section 4.2 we introduce Evaluating Levenberg-Marquardt via Kalman smoothing (ELK), a damped Newton method for numerically stable, parallel evaluation of nonlinear RNNs. We also introduce quasi-ELK, which combines quasi-DEER and ELK to create a damped Newton’s method for parallel sequential evaluation that is scalable and numerically stable.
@@ 4.1 Quasi-DEER: Scaling DEER with Diagonal Jacobian Approximations
As a consequence of our Proposition 1, replacing the Jacobians $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ with an arbitrary matrix will still result in global convergence of the resulting DEER-like algorithm in at most $T$ iterations. A straightforward way to reduce the computational cost is to replace $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ with $\left\{\mathrm{diag}\left(\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right)\right\}_{t=2}^{T}$; that is, to take the diagonal entries of the Jacobians of the dynamics functions. The parallelized version of the resulting linear recursion requires only $\mathcal{O}(DT)$ memory because it only needs to store diagonal matrices, and $\mathcal{O}(DT)$, work because the parallelized associative scan only uses elementwise vector multiplies. (Position-wise matrix-vector multiplies are still required to obtain the residuals, but this computation is easily parallelized across the sequence.) Approaches that use approximations of the Jacobian for computational reasons are referred to as quasi-Newton methods, so we refer to this algorithm as quasi-DEER. In Section 6, we show that quasi-DEER outperforms DEER in terms of wall-clock time and memory usage on the tests from Lim et al. (2024). While quasi-DEER improves the scalability of DEER, it does not address stability concerns. We propose a more stable solution below.
@@ 4.2 ELK: Stabilizing DEER with Trust Regions
Rather than treating RNN evaluation as a fixed point finding problem, let us instead consider it as an optimization problem. First, define the merit function,
@@
As in the fixed point formulation, the unique minimizer of this objective is $\mathbf{s}^{*}$. In fact, the only local minima of the merit function (7) is $\mathbf{s}^{*}$, as proved in Proposition 3 in Appendix A.2. One way of minimizing this nonlinear sum of squares objective is via the Gauss-Newton algorithm (Nocedal and Wright, 2006), which alternates between linearizing the terms in the merit function and solving the resulting linear sum-of-squares problem. The linearized objective at iteration $i$ is,
@@
The solution is $\Delta\mathbf{s}^{(i+1)}=-J(\mathbf{s}^{(i)})^{-1}\,\mathbf{r}(\mathbf{s}^{(i)})$, which is exactly the DEER update from eq. 5.
This formulation of evaluation of RNNs as nonlinear least squares also suggests more stable algorithms. For example, the Levenberg-Marquardt algorithm (Nocedal and Wright, 2006) uses updates that solve a constrained optimization problem,
@@
Finally, note that minimizing this constrained optimization problem is equivalent to minimizing a Lagrangian of the form,
@@
for some $\lambda_{i+1}\geq 0$. As noted by Särkkä and Svensson (2020), the minimizer of this Lagrangian can be obtained by a Kalman smoother. We emphasize this connection in the following proposition.
@@
Solving for the Levenberg-Marquardt update that minimizes eq. 10 with fixed $\lambda_{i+1}$ is equivalent to finding the maximum a posteriori (MAP) estimate of $\mathbf{s}_{1:T}$ in a linear Gaussian state space model, which can be done in $\mathcal{O}(\log T)$ time on a sufficiently large parallel machine.
@@
Expanding the residual and Jacobian functions in eq. 8, we see that up to an additive constant, the negative Lagrangian can be rewritten as,
^$-\widehat{\mathcal{L}}(\Delta\mathbf{s},\lambda_{i+1})\stackrel{{\scriptstyle\cdot}}{{=}}\log\mathcal{N}\left(\mathbf{s}_{1}~{}|~{}f(\mathbf{s}_{0}),I_{D}\right)+\sum_{t=1}^{T}\log\mathcal{N}\left(\mathbf{s}_{t}^{(i)}\,\Big{|}\,\mathbf{s}_{t},\frac{1}{\lambda_{i+1}}I_{D}\right)\\ +\sum_{t=2}^{T}\log\mathcal{N}\left(\mathbf{s}_{t}\,\Big{|}\,f(\mathbf{s}_{t-1}^{(i)})+\left[\frac{\partial f}{\partial\mathbf{s}}(\mathbf{s}_{t-1}^{(i)})\right](\mathbf{s}_{t-1}-\mathbf{s}_{t-1}^{(i)}),I_{D}\right),$ (11)
where $\mathcal{N}(\mathbf{x}~{}|~{}\boldsymbol{\mu},\Sigma)$ denotes the probability density function of the multivariate normal distribution.
We recognize eq. 11 as the log joint probability of a linear Gaussian state space model (LGSSM) (Särkkä, 2013) on $(\mathbf{s}_{1},\ldots,\mathbf{s}_{T})$. The means of the dynamics distributions are determined by the linearization of $f$, and the “emissions” are the states of the previous iteration, $\mathbf{s}^{(i)}$. The parameter $\lambda_{i+1}$ sets the precision of the emissions, which governs how far the posterior mode deviates from the previous states.
The minimizer of eq. 10 is equivalent to the posterior mode of the LGSSM defined in eq. 11, and can be obtained by a Kalman smoother (Särkkä, 2013). As with the linear recursions in DEER, the Kalman smoother can be implemented as a parallel associative scan that scales as in $\mathcal{O}(\log T)$ time on a machine with $\mathcal{O}(T)$ processors (Särkkä and García-Fernández, 2021; Chang et al., 2023). ∎
Therefore, we can evaluate an RNN by minimizing the merit function with the Levenberg-Marquardt algorithm. Since each step of the algorithm is performed by parallel Kalman smoothing, we call this approach Evaluating Levenberg-Marquardt via Kalman smoothing (ELK). DEER is a special case of ELK where $\lambda=0$, which can be seen as minimizing the unpenalized linearized objective eq. 8 or alternatively taking a Newton step with an infinitely large trust region. Moreover, under certain conditions, ELK also enjoys global convergence guarantees (cf. Theorems 11.7 and 11.8 of Nocedal and Wright (2006)).
@@
As with DEER, we can substitute an approximate Jacobian into the Lagrangian to obtain the quasi-ELK algorithm. Quasi-ELK enjoys the compute and memory scaling of quasi-DEER as well as stability from the trust region damping from ELK. As we show empirically in Section 6.3, while quasi-ELK takes more Newton iterates to converge than ELK, each quasi-ELK Newton iterate is faster, resulting in overall speedups in runtime.
@@
The convergence rate of ELK and quasi-ELK depends on how the trust region radius $D_{i+1}$ (or alternatively $\lambda_{i+1}$) is set. While there exist methods to analytically set $\lambda_{i}$ (Nocedal and Wright, 2006, Algorithm 4.3), these approaches require factorizing $\nicefrac{{\partial\mathbf{r}}}{{\partial\mathbf{s}}}$, which is intractable at scale. Therefore, in our experiments in Section 6.3, we simply treat $\lambda$ as a hyperparameter that can be set by a sweep over log-spaced values. We also use Kalman filtering instead of smoothing. We do so for two main reasons: filtering requires less work and memory, and we found it to converge in fewer Newton iterations than smoothing. Our current belief is that this is a result of Proposition 1, where the early part of the trace converges first, which is mirrored by filtering compared to smoothing.
@@ 5 Related Work
Nonlinear RNNs are a natural choice for modeling sequential data because of their inductive biases and memory efficiency. However, most nonlinear RNNs are not parallelizable across the sequence length. Architectures and algorithms that can exploit parallel computational hardware are core to the deep learning revolution. Therefore, a range of sequence architectures have been proposed that naturally admit parallelism, including transformers (Vaswani et al., 2017), deep linear RNNs (Martin and Cundy, 2018; Katharopoulos et al., 2020; Gu et al., 2021; Smith et al., 2023; Orvieto et al., 2023; Gu and Dao, 2023; Beck et al., 2024), and convolution approaches (Oord et al., 2016; Romero et al., 2021; Poli et al., 2023; Parnichkun et al., 2024). These approaches obtain parallelization by developing new architectures, and do not consider parallel evaluation of existing nonlinear RNN architectures. DEER (Lim et al., 2024) is notable as it considers parallel evaluation and training of arbitrary nonlinear RNNs.
@@
Beyond DEER, there has been broad interest in using root finders/fixed-point iterations in deep learning and sequence modeling. Deep implicit layers (Kolter et al., 2020; Bai et al., 2019, 2020, 2021) and neural ODEs (Chen et al., 2018; Massaroli et al., 2021) replace conventional feed forward network layers (Goodfellow et al., 2016) with an implicit layer whose output is the root of an equation. Parallel decoding methods (Mihaylova and Martins, 2019; Santilli et al., 2023; Fu et al., 2024) adapt ideas from Jacobi and Gauss-Siedel iterations to evaluate autoregressive sequence models in parallel. However, these methods presuppose the existence of a parallelized forward pass for the sequence model and do not leverage additional gradient information to obtain sublinear convergence.
@@
There is a large literature on scaling Newton’s method and making it globally convergent. Quasi-Newton methods are efficient algorithms that use an approximation of the Jacobian or Hessian in Newton’s method, and include approaches like BFGS (Broyden, 1970; Fletcher, 1970; Goldfarb, 1970; Shanno, 1970) and L-BFGS (Liu and Nocedal, 1989). Other approaches use Newton’s method to optimize deep nets (Martens and Grosse, 2015). However, these quasi-Newton algorithms do not admit efficient parallel scans. There are also conjugate gradients methods for exploiting structured Jacobians or Hessians (Steihaug, 1983), though they often do not attain the fast convergence rates of Newton or quasi-Newton methods (Nocedal and Wright, 2006). Methods for stabilizing and ensuring Newtons method converges globally include regularization approaches (Nesterov and Polyak, 2006; Doikov and Nesterov, 2023), backtracking line search (Ortega and Rheinboldt, 2000), and trust regions (Conn et al., 2000). All these stabilization methods have strengths and weaknesses, but as noted by Nocedal and Wright (2006), “the trust-region Newton method has proved to be highly effective in practice,” leading us to apply it to evaluating RNNs.
@@
Bell and Cathey (1993) and Bell (1994) draw connections between the Gauss-Newton method and the iterated extended Kalman filter and smoother (Sorenson, 1966; Särkkä, 2013). Because Gauss-Newton is unstable, it is natural to use Levenberg-Marquardt (Levenberg, 1944; Marquardt, 1963) to stabilize the filtering/smoothing problem (Chen and Oliver, 2013; Mandel et al., 2016; Särkkä and Svensson, 2020). These approaches start with a smoothing problem, and stabilize it using approaches from nonlinear equations; whereas we start with a nonlinear equation to solve, and make the connection with Kalman filtering to leverage parallelized algorithms (Särkkä and García-Fernández, 2021). We also address the practicalities of operationalizing this connection for modern deep networks.
@@ 6 Experiments
We now experimentally examine the relative performance of sequential evaluation, and the full- and quasi- variants of DEER and ELK. Specifically, we examine the following:
  • Can quasi-DEER provide memory savings over DEER and runtime savings over sequential evaluation, while retaining the accuracy of training and evaluation?
  • Can ELK and quasi-ELK be used to stabilize evaluation in regimes where DEER is unstable; and what are the relative strengths and weaknesses of each variant?
We tackle the first point in Sections 6.1 and 6.2. We replicate experiments from Lim et al. (2024) and show quasi-DEER retains the parallelized runtime and accuracy of DEER, and can reduce memory consumption by up to an order of magnitude. In Section 6.3 we show that ELK and quasi-ELK remain stable in a domain whereas DEER becomes unstable, with quasi-ELK being the fastest of all parallelized methods. We provide further experimental details in Appendix B.

### v3

@@ 4 Scaling and Stabilizing Newton’s Method for Parallel Evaluation
In Section 4.1 we introduce quasi-DEER, a quasi-Newton method that addresses the intractability of DEER for large state sizes. In Section 4.2 we introduce Evaluating Levenberg-Marquardt with Kalman (ELK), a damped Newton method for numerically stable, parallel evaluation of nonlinear RNNs. We also introduce quasi-ELK, which combines quasi-DEER and ELK to create a damped Newton’s method for parallel sequential evaluation that is scalable and numerically stable.
@@ 4.1 Quasi-DEER: Scaling DEER with Diagonal Jacobian Approximations
As a consequence of our Proposition 1, replacing the Jacobians $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ with an arbitrary matrix will still result in global convergence of the resulting DEER-like algorithm in at most $T$ iterations. A straightforward way to reduce the computational cost is to replace $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ with $\left\{\mathrm{diag}\left(\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right)\right\}_{t=2}^{T}$, i.e. take the diagonal entries of the Jacobians of the dynamics functions. The resulting linear recursion requires only $\mathcal{O}(TD)$ memory because it only needs to store diagonal matrices, and $\mathcal{O}(TD)$ work, because the parallelized associative scan only uses element-wise vector multiplies. Position-wise matrix-vector multiplies are still required to obtain the residuals, but this computation can be naively parallelized across the sequence.
Quasi-Newton methods approximate the Jacobian for computational reasons, so we refer to this algorithm as quasi-DEER. In Section 6, we show that quasi-DEER outperforms DEER on wall-clock time and memory usage on the tests from Lim et al. [1]. Quasi-DEER improves the scalability of DEER, but it does not address stability concerns. We propose a more stable solution below.
@@ 4.2 ELK: Stabilizing DEER with Trust Regions
Rather than treating RNN evaluation as a fixed-point finding problem, let us instead consider it as an optimization problem. First, we define the merit function
@@
As in the fixed-point formulation, the unique minimizer of this objective is $\mathbf{s}^{*}$. In fact, the only local minimum of the merit function (7) is $\mathbf{s}^{*}$, as proved in Proposition 3 in Appendix A.2. One way of minimizing this nonlinear sum of squares objective is via the Gauss-Newton algorithm [23], which alternates between linearizing the terms in the merit function and solving the resulting linear sum-of-squares problem. The linearized objective at iteration $i$ is
@@
The solution is $\Delta\mathbf{s}^{(i+1)}=-J(\mathbf{s}^{(i)})^{-1}\,\mathbf{r}(\mathbf{s}^{(i)})$, which is exactly the DEER update from (5).
Formulating evaluation as nonlinear least squares also suggests more stable algorithms. The Levenberg-Marquardt algorithm [23] uses updates that solve a constrained optimization problem
@@
Finally, minimizing this constrained optimization is equivalent to minimizing the Lagrangian
@@
for some $\lambda_{i+1}\geq 0$. As noted by Särkkä and Svensson [25], the minimizer of this Lagrangian can be obtained by a Kalman smoother. We emphasize this connection in the following proposition.
@@
Solving for the Levenberg-Marquardt update that minimizes (10) with fixed $\lambda_{i+1}$ is equivalent to finding the maximum a posteriori (MAP) estimate of $\mathbf{s}_{1:T}$ in a linear Gaussian state space model, which can be done in $\mathcal{O}(\log T)$ time on a sufficiently large parallel machine.
@@
Expanding the residual and Jacobian functions in (8), we see that up to an additive constant, the negative Lagrangian can be rewritten as,
^$-\widehat{\mathcal{L}}(\Delta\mathbf{s},\lambda_{i+1})\stackrel{{\scriptstyle\cdot}}{{=}}\log\mathcal{N}\left(\mathbf{s}_{1}\mid f(\mathbf{s}_{0}),I_{D}\right)+\sum_{t=1}^{T}\log\mathcal{N}\left(\mathbf{s}_{t}^{(i)}\,\Big{|}\,\mathbf{s}_{t},\frac{1}{\lambda_{i+1}}I_{D}\right)\\ +\sum_{t=2}^{T}\log\mathcal{N}\left(\mathbf{s}_{t}\,\Big{|}\,f(\mathbf{s}_{t-1}^{(i)})+\left[\frac{\partial f}{\partial\mathbf{s}}(\mathbf{s}_{t-1}^{(i)})\right](\mathbf{s}_{t-1}-\mathbf{s}_{t-1}^{(i)}),I_{D}\right),$ (11)
where $\mathcal{N}(\mathbf{x}\mid\bm{\mu},\Sigma)$ denotes the probability density function of the multivariate normal distribution.
We recognize (11) as the log joint probability of a linear Gaussian state space model (LGSSM) [31] on $(\mathbf{s}_{1},\ldots,\mathbf{s}_{T})$. The means of the dynamics distributions are given by the linearization of $f$, and the emissions are the previous iteration’s states, $\mathbf{s}^{(i)}$. The parameter $\lambda_{i+1}$ sets the precision of the emissions, governing how far the posterior mode deviates from the previous states.
The minimizer of (10) is the posterior mode of the LGSSM (11), and can be obtained by Kalman smoothing [31]. As with the linear recursions in DEER, the Kalman smoother can be implemented as a parallel scan that scales as $\mathcal{O}(\log T)$ in time on a machine with $\mathcal{O}(T)$ processors [26, 27]. ∎
Therefore, we can evaluate an RNN by minimizing the merit function with the Levenberg-Marquardt algorithm. Since each step of the algorithm is performed by parallel Kalman smoothing, we call this approach Evaluating Levenberg-Marquardt with Kalman (ELK). Note that DEER is a special case of ELK, where $\lambda=0$, which can be seen as minimizing the unpenalized linearized objective (8), or, alternatively, taking a Newton step with an infinitely large trust region. Moreover, under certain conditions, ELK also enjoys global convergence guarantees [23, Thms. 11.7, 11.8].
@@
As with DEER, we can substitute an approximate Jacobian into the Lagrangian to obtain the quasi-ELK algorithm. Quasi-ELK enjoys the compute and memory scaling of quasi-DEER, as well as stability from the trust region damping from ELK. We show empirically in Section 6.3 that while quasi-ELK takes more iterates to converge than ELK, each quasi-ELK iterate is faster, giving overall runtime speedups.
@@
The convergence rate of (quasi-)ELK depends on the trust region radius $D_{i}$ (or alternatively $\lambda_{i}$). Although there exist methods to analytically set $\lambda_{i}$ [23, Algorithm 4.3], these approaches require factorizing $\nicefrac{{\partial\mathbf{r}}}{{\partial\mathbf{s}}}$, which is intractable at scale. Therefore, we treat $\lambda$ as a hyperparameter set by a sweep over log-spaced values (cf. Appendix B.4).
We also use Kalman filtering instead of smoothing. We do so for two main reasons: filtering requires less work and memory; and we also found it to converge in fewer Newton iterations than smoothing. We hypothesize that this follows from Proposition 1, where the early part of the trace converges first. In Appendix A.3 we also discuss an alternative interpretation of ELK and the Kalman filter as defining a linear recurrence where the trust region attenuates the eigenvalues used in the parallel scan.
Limitations
The quasi-Newton methods lose the local quadratic convergence properties of Newton but remain globally convergent (cf. Proposition 1). Our implementation of quasi-DEER for training uses approximate gradients (cf. Section 6.2). The heuristic of resetting to the states to zero when they become unstable is also motivated by Proposition 1, but it slows convergence in (quasi-)DEER methods. As a result, we develop ELK to stabilize evaluation. Like DEER, ELK has cubic complexity in $D$, which we combat with quasi-ELK. However, quasi-ELK adds an additional hyperparameter. Note that all four parallelized methods discussed in this paper, as well as sequential evaluation of RNNs, have different regimes where they are fastest. For example, in our evaluation of autoregressive RNNs (Section 6.3), the ELK methods are faster than the DEER methods on wall-clock time, but they are slower than sequential. In our evaluation of the Lorenz96 system (Appendix B.5), ELK is more stable than DEER, but DEER is faster on wall-clock time. An area for future research is characterizing the properties of dynamical systems and hardware where each method is fastest. Finally, at the core of the implementation of the parallelized methods is the parallel associative scan (cf. Appendix B.6), which is currently available in JAX [30] but not in the standard PyTorch package.
@@ 5 Related Work
Nonlinear RNNs are a natural choice for modeling sequential data because of their inductive biases and memory efficiency. However, most nonlinear RNNs are not parallelizable over the sequence length, and architectures that can exploit parallel computational hardware have been core to the success of deep learning. Therefore, a range of sequence architectures that inherently admit parallelism have been proposed, including transformers [3], deep linear RNNs [4, 5, 6, 7, 8, 32, 33, 34, 35], and convolutions [36, 37, 38, 39]. These methods obtain parallelism by developing new architectures, and do not consider parallelizing existing nonlinear architectures. DEER [1] is notable as it considers parallel evaluation and training of arbitrary nonlinear RNNs.
@@
Beyond DEER, there has been broad interest in using root finders/fixed-point iterations in deep learning and sequence modeling. Deep implicit layers [40, 41, 42, 43] and neural ODEs [28, 44] replace conventional feed forward network layers [45] with an implicit layer whose output is the root of an equation. Moreover, Song et al. [46] parallelize the evaluation of feedforward nets using Jacobi and Gauss-Siedel iterations. In sequence modeling, parallel decoding methods [47, 48, 49] adapt ideas from Jacobi and Gauss-Siedel iterations to evaluate autoregressive sequence models in parallel. These approaches iteratively refine inputs by repeatedly feeding in previous outputs back into a parallelized sequence model. However, these methods presuppose the existence of a parallelized forward pass for the sequence model and do not leverage additional gradient information to obtain sublinear convergence.
Parallelizing Dynamical Systems over Time
Other work has investigated evaluating other nonlinear dynamical systems over time. ParaDIGMS [50] parallelizes sampling from diffusion models but uses Picard iterations instead of Newton’s method, while Selvam et al. [51] use Parareal iterations. In the numerical ODE and PDE communities there has been great interest in Parallel in Time methods; see Gander [52], Ong and Schroder [53] for surveys. Vargas et al. [54] parallelized the evaluation of chaotic dynamical systems over time, but instead of casting Newton’s method as a parallel scan, they resort to multi-grid methods to evaluate at different hierarchies. Moreover, these methods have not yet been applied to parallelizing RNNs.
@@
Quasi-Newton methods are efficient algorithms that use an approximation of the Jacobian or Hessian in Newton’s method, and include approaches like BFGS [55, 56, 57, 58] and L-BFGS [59]. Other approaches use Newton’s method to optimize deep nets [60]. However, these quasi-Newton algorithms do not admit efficient parallel scans. There are also conjugate gradients methods for exploiting structured Jacobians or Hessians [61], though they often do not attain the fast convergence rates of Newton or quasi-Newton methods [23]. Methods for stabilizing and ensuring Newton’s method converges globally include regularization approaches [62, 63], backtracking line search [64], and trust regions [65]. All these stabilization methods have strengths and weaknesses, but as noted by Nocedal and Wright [23]: “the trust-region Newton method has proved to be highly effective in practice,” leading us to apply it to evaluating RNNs.
@@
Bell and Cathey [66] and Bell [67] draw connections between the Gauss-Newton method and the iterated extended Kalman filter and smoother [68, 31]. Because Gauss-Newton is unstable, it is natural to use Levenberg-Marquardt [69, 70] to stabilize the filtering/smoothing problem [25, 71, 72]. These approaches start with a smoothing problem and stabilize it using approaches from nonlinear equations, whereas we start with a nonlinear equation to solve and make the connection with Kalman filtering to leverage parallelized algorithms [26]. We also discuss the practicalities of this connection for modern deep networks.
@@ 6 Experiments
We now experimentally examine the relative performance of these methods. Specifically, we evaluate: 1) whether quasi-DEER can provide memory savings over DEER and runtime savings over sequential evaluation, while retaining the accuracy of training and evaluation; and 2) whether ELK and quasi-ELK can be used to stabilize evaluation in regimes where DEER is unstable. In Sections 6.1 and 6.2 we use experimental designs from Lim et al. [1] and show that quasi-DEER retains the fast runtime and accuracy of DEER and can reduce memory consumption by up to an order of magnitude. In Section 6.3 we show that (quasi-)ELK remains stable when DEER becomes unstable, and that quasi-ELK is the fastest of all parallelized methods. We provide further details in Appendix B.